# Primus: A Pioneering Collection of Open-Source Datasets for Cybersecurity LLM Training

### v1

@@ 2.5 Primus-Reasoning
Table 4: Statistics of the Primus-Reasoning dataset, distilled from o1-preview using CTI-Bench, with only accepted correct samples.
@@
3 Evaluation Protocol
In this section, we first introduce the cybersecurity benchmarks used to evaluate training performance (Sec.3.1), followed by the specific evaluation settings (Sec.3.2).
3.1 Benchmarks
CISSP.
The Certified Information Systems Security Professional (CISSP) is a widely recognized certification in the field of cybersecurity. It assesses both technical expertise and managerial competence in designing, building, and managing an organization’s security posture. We construct an evaluation set based on multiple-choice questions taken from the assessment tests within the CISSP learning materials.
CTI-Bench.
As introduced in Sec.2.5, CTI-Bench is a benchmark for evaluating the reasoning and knowledge capabilities of LLMs in CTI. It consists of several subtasks, including CTI-RCM, CTI-VSP, CTI-ATE, and CTI-MCQ, which assess a model’s ability to analyze vulnerabilities, infer security risks, extract attack techniques, and understand cybersecurity concepts.
CyberMetric.
CyberMetric Tihanyi et al. (2024) is a widely recognized benchmark designed to assess LLMs’ cybersecurity knowledge across multiple domains. It includes high-quality, human-verified multiple-choice questions covering cryptography, network security, penetration testing, and compliance. We select the 500-question subset for evaluation as it provides a balanced and representative assessment of cybersecurity knowledge.
SecEval.
SecEval Li et al. (2023a) is a benchmark specifically designed to assess cybersecurity knowledge. It consists of over 2,000 multiple-choice questions across nine domains, including software security, cryptography, and network security. The dataset was generated by prompting GPT-4 with authoritative sources such as open-licensed textbooks, official documentation, and industry standards. Given its broad coverage and rigorous quality control, SecEval serves as a reliable benchmark for assessing the cybersecurity proficiency of LLMs.
3.2 Evaluation Settings
5-shot, w/o Chain-of-Thought (CoT).
0-shot, w/ CoT.
We follow the evaluation setup used in the OpenAI technical report benchmarks with simple-eval [^10] https://github.com/openai/simple-evals, using a standardized prompt [^11] The prompt is provided in the Appx.A (Fig.11) that allows the model to articulate its reasoning before producing the final answer. Due to the formatting variability of CoT responses, we use GPT-4o-mini to extract the final answers before scoring.
4 Training and Results
4.1 Overview
In this section, we present the entire training pipeline, which consists of four key stages. First, we expand the model’s cybersecurity expertise and understanding through continual pre-training (Sec.4.2), which reinforces key cybersecurity concepts and enables the model to provide accurate information on security threats and mitigation strategies. Next, we restore its instruction-following capability through instruction fine-tuning (Sec.4.3), and further refine it through model merging to balance instruction-following and cybersecurity expertise. Finally, we train the model to develop reasoning capabilities on cybersecurity tasks (Sec.4.4) [^12] The training hyperparameters for each stage are provided in the Appx.B.
4.2 Pre-Training
+ Primus-Seed 0.7132 0.6608 0.6100 1.2848 0.2829 0.8600 0.4998 2.34$\uparrow$2.10%
+ Primus-FineWeb 0.7191 0.6600 0.6680 1.1499 0.3006 0.8620 0.4984 2.56$\uparrow$11.53%
+ Primus-Seed+FineWeb 0.7230 0.6676 0.6780 1.0912 0.3140 0.8660 0.5007 2.66$\uparrow$15.88%
Table 5: Performance of continual pretraining on Llama across cybersecurity benchmarks. The last three rows indicate pretraining with Primus-Seed, Primus-FineWeb, and their combination. CTI-VSP is scored using Mean Absolute Deviation (lower is better), CTI-ATE uses F1 score, and the others use accuracy. The aggregate score (Agg.) is the sum of all benchmarks, with CTI-VSP negated. The best results are highlighted in bold.
We use Llama-3.1-8B-Instruct as our base model due to its wide community adoption and strong performance at the same parameter scale. We perform continual pre-training on two cybersecurity datasets: Primus-Seed (Sec.2.2), which consists of curated cybersecurity text, and Primus-FineWeb (Sec.2.3), a filtered subset of cybersecurity content from FineWeb, to expand the model’s cybersecurity expertise and understanding. To assess performance improvements, we evaluate the model against the seven cybersecurity benchmarks described in Sec.3.1 (5-shot, w/o CoT).
We train the model using the NeMo NVIDIA (2025) on four 8$\times$×H200 nodes, with training hyperparameters and details provided in Appx.B. To analyze the impact of different datasets, we conduct an ablation study by pre-training the model separately on each dataset and jointly on both for two epochs. The results in Tab.5 show that pre-training on either dataset improves the cybersecurity performance in the aggregate evaluation score. However, the largest improvement, 15.88%, is observed when pre-training on the combined dataset, so we adopt this model as the Llama-Primus-Base for subsequent training stages.
4.3 Instruction Fine-Tuning and Merge
Llama-Primus-Instruct maintains its advantage in cybersecurity while achieving an MT-Bench score of 7.91. However, this remains lower than the 8.35 of Llama, resulting in a limited improvement in the aggregated score (2.4%). To mitigate this, we apply DARE-TIES Yu et al. (2024); Yadav et al. (2023), a model merging technique that balances diverse capabilities—specifically, instruction-following and cybersecurity expertise in our case. We conduct a grid search over the merging ratio, setting Llama-Primus-Instruct:Llama-3.1-8B-Instruct to $(0.5+w)$:$(0.5-w)$ and varying $w$ from 0 to 0.5 in steps of 0.05. The optimal ratio that maximizes the aggregated score is found to be 0.75:0.25, with the merged model chosen as Llama-Primus-Merged. Notably, this configuration retains cybersecurity performance comparable to Llama-Primus-Instruct while restoring the MT-Bench to 8.29, almost equal to Llama, resulting in a 5.4% improvement in the aggregated score.
4.4 Reasoning Fine-Tuning
Model CISSP Avg. Tokens
w/o CoT, 5-shot
Llama-3.1-8B-Instruct 0.7073 1
Llama-Primus-Merged 0.7191 $\uparrow$1.671
w/ CoT, 0-shot
Llama-3.1-8B-Instruct 0.7288$\uparrow$3.03279.69
DeepSeek-R1-Distill-Llama-8B 0.7399$\uparrow$4.611542.10
Llama-Primus-Merged 0.7603$\uparrow$7.49241.92
Finetuned on Primus-Reasoning
Llama-3.1-8B-Reasoning 0.7583$\uparrow$7.21646.94
Llama-Primus-Reasoning 0.7780$\uparrow$10.0726.96
o1-preview 0.8035 1054.91
Table 7: Effect of Primus-Reasoning fine-tuning, evaluated on CISSP. $\uparrow$ indicates the percentage improvement over Llama without CoT and in the 5-shot setting. The best improvement is highlighted in bold.
We further distill Llama-Primus-Merged using Primus-Reasoning (Sec.2.5), a high-quality dataset of cybersecurity task reasoning steps obtained from o1-preview, to equip it with reasoning and self-reflection capabilities. This approach has been successfully demonstrated in previous work such as S1 Muennighoff et al. (2025) and Sky-T1 Team (2025). Since Primus-Reasoning is constructed from CTI-Bench tasks, we exclude them from the evaluation and choose CISSP as a representative metric, as it also emphasizes reasoning rather than just factual recall. The results are presented in Tab.7.
As shown in the table, both Llama-3.1-8B-Instruct and Llama-Primus-Merged improve with CoT over direct answer generation. Notably, Llama-Primus-Merged achieves the largest gain, even outperforming DeepSeek-R1-Distill-Llama-8B [^13] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B with the fewest tokens used, suggesting that stronger cybersecurity knowledge benefits reasoning. After fine-tuning on Primus-Reasoning (lower part of the table), we observe that reasoning tokens usage triples while accuracy improves further, with Llama-Primus-Reasoning achieving the largest improvement (10%). Interestingly, comparing Llama-3.1-8B-Reasoning and DeepSeek-R1-Distill-Llama-8B may suggest that domain-specific reasoning distillation yields better in-domain performance than general-domain distillation.
5 Conclusion
In this work, we explore the adaptation of other successful domain-specific LLM approaches to cybersecurity and contribute a series of datasets covering different stages of LLM training, including pre-training, instruction fine-tuning, and reasoning distillation, each of which has been validated to improve cybersecurity performance. To our knowledge, this is the first study to systematically strengthen the cybersecurity skills of an LLM across multiple stages of training, and we will release all datasets and models to encourage further community research.
Limitations
• Due to limited computational resources, our experiments are limited to 8B scale models, leaving the effectiveness of scaling to larger models (e.g., 70B or 405B) unknown.
Ethics Statement
We used Garak Derczynski et al. (2024), a toolkit that probes for hallucination, data leakage, prompt injection, misinformation, toxicity generation, jailbreaks, and many other vulnerabilities, to evaluate Llama-Primus-Merged. The results showed no significant differences compared to Llama (Appx.C). However, we still emphasize that the user is are solely responsible for the content generated with the Primus model, as it lacks mechanisms to handle the disclosure of harmful, biased, or toxic content. Therefore, we strongly recommend that Primus be used for research purposes only. If used in production for natural language generation, users should independently assess the risks and implement appropriate safeguards.
Appendix A Prompts

### v2

@@ 2.5 Primus-Reasoning
Table 4: Statistics of the Primus-Reasoning dataset, distilled from o1-preview and DeepSeek-R1 on CTI-Bench questions, with only accepted correct samples.
3 Evaluation Protocol
This section introduces the cybersecurity benchmarks (Sec.3.1) and evaluation settings (Sec.3.2) used to assess training performance.
3.1 Benchmarks
CISSP.
The Certified Information Systems Security Professional (CISSP) is a widely recognized cybersecurity certification that assesses both technical expertise and managerial competence. We construct an evaluation set based on multiple-choice questions from CISSP learning materials.
CTI-Bench.
CTI-Bench is a benchmark for evaluating the reasoning and knowledge capabilities of LLMs in CTI. It consists of several subtasks, including CTI-RCM, CTI-VSP, CTI-ATE, and CTI-MCQ, which assess a model’s ability to analyze vulnerabilities, infer security risks, extract attack techniques, and understand cybersecurity concepts.
CyberMetric.
CyberMetric Tihanyi et al. (2024) is a benchmark of human-verified multiple-choice questions designed to assess LLMs’ cybersecurity knowledge across domains such as cryptography, network security, penetration testing, and compliance. We select a 500-question subset for evaluation as it is balanced and representative.
SecEval.
SecEval Li et al. (2023a) is a benchmark consisting of over 2,000 multiple-choice questions covering nine cybersecurity domains, including software security, cryptography, and network security. Generated by prompting GPT-4 with authoritative sources such as textbooks and official documentation, it provides a reliable measure of LLMs’ cybersecurity proficiency.
3.2 Evaluation Settings
5-shot, w/o Chain-of-Thought (CoT).
0-shot, w/ CoT.
We follow the evaluation setup from the OpenAI technical report benchmarks with simple-eval [^11] https://github.com/openai/simple-evals, using a standardized prompt [^12] The prompt is provided in the Appx.E (Fig.13) that allows the model to articulate its reasoning before producing the final answer. Due to the formatting variability of CoT responses, we use GPT-4o-mini to extract the final answers before scoring.
4 Training and Results
4.1 Overview
In this section, we present the entire training pipeline, which consists of four key stages. First, we expand the model’s cybersecurity expertise and understanding through continued pre-training (Sec.4.2), which reinforces key cybersecurity concepts and enables the model to provide accurate information on security threats and mitigation strategies. Next, we restore its instruction-following capability through instruction fine-tuning (Sec.4.3), and further refine it through model merging to balance instruction-following and cybersecurity expertise. Finally, we train the model to develop reasoning capabilities on cybersecurity tasks (Sec.4.4) [^13] The training hyperparameters for each stage are provided in the Appx.D.
4.2 Pre-Training
+ Primus-Seed 0.7132 0.6608 0.6100 1.2848 0.2829 0.8600 0.4998 2.34$\uparrow$2.1%
+ Primus-FineWeb 0.7191 0.6600 0.6680 1.1499 0.3006 0.8620 0.4984 2.56$\uparrow$11.5%
+ Primus-Seed+FineWeb 0.7230 0.6676 0.6780 1.0912 0.3140 0.8660 0.5007 2.66$\uparrow$15.9%
Table 5: Performance of continued pretraining on Llama across cybersecurity benchmarks. The last three rows indicate pretraining with Primus-Seed, Primus-FineWeb, and their combination. CTI-VSP is scored using Mean Absolute Deviation (lower is better), CTI-ATE uses F1 score, and the others use accuracy. The aggregate score (Agg.) is the sum of all benchmarks, with CTI-VSP negated. The best results are highlighted in bold.
We use Llama-3.1-8B-Instruct as our base model due to its wide community adoption and strong performance at the same parameter scale. We perform continued pre-training on two cybersecurity datasets: Primus-Seed (Sec.2.2), which consists of curated cybersecurity text, and Primus-FineWeb (Sec.2.3), a filtered subset of cybersecurity content from FineWeb, to expand the model’s cybersecurity expertise and understanding. To assess performance improvements, we evaluate the model against the seven cybersecurity benchmarks described in Sec.3.1 (5-shot, w/o CoT).
We train the model using the NeMo NVIDIA (2025) on four 8$\times$×H200 nodes, with training hyperparameters and details provided in Appx.D. To analyze the impact of different datasets, we conduct an ablation study by pre-training the model separately on each dataset and jointly on both for two epochs. The results in Tab.5 show that pre-training on either dataset improves the cybersecurity performance in the aggregate evaluation score. However, the largest improvement, 15.9%, is observed when pre-training on the combined dataset, so we adopt this model as the Llama-Primus-Base for subsequent training stages [^14] We also experimented with a 70B model in Q2 of Appx.A (FAQs).
4.3 Instruction Fine-Tuning and Merge
Llama-Primus-Instruct maintains its advantage in cybersecurity while achieving an MT-Bench score of 7.91. However, this remains lower than the 8.35 of Llama, resulting in a limited improvement in the aggregated score (2.4%). To mitigate this, we apply DARE-TIES Yu et al. (2024); Yadav et al. (2023), a model merging technique that balances diverse capabilities—specifically, instruction-following and cybersecurity expertise in our case. We conduct a grid search over the merging ratio, setting Llama-Primus-Instruct:Llama-3.1-8B-Instruct to $(0.5+w)$:$(0.5-w)$ and varying $w$ from 0 to 0.5 in steps of 0.05. The optimal ratio that maximizes the aggregated score is found to be 0.75:0.25, with the merged model chosen as Llama-Primus-Merged. Notably, this configuration retains cybersecurity performance comparable to Llama-Primus-Instruct while restoring the MT-Bench to 8.29, almost equal to Llama, resulting in a 5.4% improvement in the aggregated score [^15] We provide more details in Q4 and Q5 of Appx.A (FAQs).
4.4 Reasoning Fine-Tuning
Model CISSP Avg. Tokens
w/o CoT, 5-shot
Llama-3.1-8B-Instruct 0.7073 1
Llama-Primus-Merged 0.7191 $\uparrow$1.671
w/ CoT, 0-shot
Llama-3.1-8B-Instruct 0.7288 $\uparrow$3.03279.69
  + Distilled from o1-preview 0.7583 $\uparrow$7.21646.94
  + Distilled from DeepSeek-R1 0.7859 $\uparrow$11.11667.56
  + Distilled from (o1 + R1) 0.7780 $\uparrow$10.01615.54
Llama-Primus-Merged 0.7603 $\uparrow$7.49241.92
  + Distilled from o1-preview 0.7780 $\uparrow$10.0726.96
  + Distilled from DeepSeek-R1 0.8075 $\uparrow$14.21483.94
  + Distilled from (o1 + R1) 0.8193 $\uparrow$15.81467.40
o1-preview 0.8035 1054.91
DeepSeek-R1 0.8212 1229.32
DeepSeek-R1-Distill-Llama-8B 0.7399 1542.10
Table 7: Effect of Primus-Reasoning fine-tuning (on o1-preview, DeepSeek-R1, and their combination), evaluated on CISSP. $\uparrow$ indicates the percentage improvement over Llama without CoT and in the 5-shot setting. The best improvement is highlighted in bold.
We further distill Llama-Primus-Merged using Primus-Reasoning (Sec.2.5), a high-quality dataset of cybersecurity task reasoning steps obtained from o1-preview and DeepSeek-R1, to equip it with reasoning and self-reflection capabilities. This approach has been successfully demonstrated in previous work such as S1 Muennighoff et al. (2025) and Sky-T1 Team (2025). Since Primus-Reasoning is constructed from CTI-Bench tasks, we exclude them from the evaluation and choose CISSP as a representative metric, as it also emphasizes reasoning rather than just factual recall. The results are presented in Tab.7.
As shown in the table, both Llama-3.1-8B-Instruct and Llama-Primus-Merged improve with CoT over direct answer generation. Notably, Llama-Primus-Merged achieves the largest gain, even outperforming DeepSeek-R1-Distill-Llama-8B [^16] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B (0.7603 vs. 0.7399) with the fewest tokens,suggesting stronger cybersecurity knowledge benefits reasoning. After fine-tuning on Primus-Reasoning (rows starting with +), token usage increases while accuracy further improves; distillation on the combined o1-preview and DeepSeek-R1 data achieves the largest improvement (15.8%). Interestingly, comparing DeepSeek-R1-Distill-Llama-8B (0.7399) and Llama-3.1-8B-Instruct after distillation (0.7583 / 0.7859 / 0.7780) may suggest that domain-specific reasoning distillation yields better in-domain performance than general-domain distillation.
5 Domain Calibration Analysis
Benchmark ECE (%)
Llama-3.1- 8B-Instruct Llama-Primus- Base Llama-Primus- Merged
CISSP 7.22 4.59 4.55
CTI-MCQ 11.01 2.03 5.52
CyberMetric 4.11 3.41 2.57
Average 7.45 3.34$\downarrow$55.174.21$\downarrow$43.49%
Table 8: Expected Calibration Error (ECE) across cybersecurity benchmarks (with 10 bins).
Metric Llama-3.1- 8B-Instruct Llama-Primus- Base Llama-Primus- Merged
Accuracy (%) 67.56 66.29 66.59
ECE (%) 5.99 6.07 5.56
Table 9: Accuracy and ECE across models on MMLU.
In cybersecurity applications, a model’s confidence score is often a critical indicator for deciding whether to escalate issues for human intervention, such as sending alerts to security analysts. For this to work, the confidence score must accurately reflect the true accuracy. After multi-stage training in the cybersecurity domain, we found that our model had a significantly lower Expected Calibration Error (ECE) Guo et al. (2017) on cybersecurity-related questions. This suggests our model’s confidence is more aligned with its actual accuracy. The ECE measures the average discrepancy between a model’s confidence and its empirical accuracy.
Specifically, we re-evaluated the cybersecurity multiple-choice tasks (CISSP, CTI-MCQ, and CyberMetric). We took the token probability of the output answer (A/B/C/D) as the confidence score and calculated the ECE, as shown in Tab.8. The ECE of our model on cybersecurity questions was reduced by half, indicating that the model is better calibrated and thus more reliable in practical applications, especially those involving confidence thresholds. Additionally, evaluation on general-domain questions (e.g., MMLU) Hendrycks et al. (2021) showed no significant change (see Tab.9).
Recent work has sought to improve LLM calibration by reducing ECE through specialized training methods Xu et al. (2024b). However, leveraging domain-specific data for this purpose remains unexplored. We posit that our approach could provide valuable insights into confidence calibration.
6 Conclusion
In this work, we explore adapting other successful domain-specific LLM approaches to cybersecurity and contribute a series of datasets covering different stages of LLM training, including pre-training, instruction fine-tuning, and reasoning distillation, each of which has been validated to improve cybersecurity performance. To our knowledge, this is the first study to systematically strengthen the cybersecurity skills of an LLM across multiple stages of training, and we will release all datasets and models to encourage further community research.
Limitations
• Due to limited computational resources, our experiments primarily focus on 8B-scale models, leaving the effectiveness of scaling to larger models (e.g., 405B or 671B) unknown.
Ethics Statement
We used Garak Derczynski et al. (2024), a toolkit that probes for hallucination, data leakage, prompt injection, misinformation, toxicity generation, jailbreaks, and many other vulnerabilities, to evaluate Llama-Primus-Merged. The results showed no significant differences compared to Llama (Appx.F). However, we still emphasize that the user is solely responsible for the content generated with the Primus model, as it lacks mechanisms to handle the disclosure of harmful, biased, or toxic content. Therefore, we strongly recommend that Primus be used for research purposes only. If used in production for natural language generation, users should independently assess the risks and implement appropriate safeguards.
Appendix A FAQs
• Q1: What are the implementation details, such as the training hyperparameters and the prompts used for the LLM during dataset construction?
These details are provided in the appendix. The training hyperparameters are listed in Appx.D, and the prompts used for dataset construction are included in Appx.E.
• Q2: The experiments primarily target 8B models. A natural follow-up is whether these datasets generalize to larger models, i.e., whether they can also improve the cybersecurity performance of larger models?
Llama-3.1-Nemotron-70B-Instruct 0.8527 0.6900 0.6590 1.1893 0.3905 0.9380 0.7177 3.06
Llama-Primus-Nemotron-70B-Base 0.8703 0.7148 0.7410 1.0281 0.4540 0.9280 0.7208 3.40$\uparrow$11.2%
Table 10: Performance comparison of Llama-3.1-Nemotron-70B-Instruct and Llama-Primus-Nemotron-70B-Base on cybersecurity benchmarks. CTI-VSP is scored using Mean Absolute Deviation (lower is better), CTI-ATE uses F1 score, and the others use accuracy. The aggregate score (Agg.) is the sum of all benchmarks, with CTI-VSP negated. The best results are highlighted in bold.
Yes, we extended our experiments to a 70B model by further pre-training Llama-3.1-Nemotron-70B-Instruct to obtain Llama-Primus-Nemotron-70B-Base. In addition to the dataset used for the 8B model, we supplemented its pre-training corpus with 7.6B tokens of cybersecurity content filtered from Nemotron-CC Su et al. (2024) (see Appx.B). The results in Tab.10 show an 11.2% gain in the aggregated cybersecurity benchmark score. We will also release this model under the MIT license. Due to its high computational cost, we did not conduct the dataset-combination ablation study on the 70B model that we performed on the 8B experiments.
• Q3: Since LLMs (e.g., Claude) were used during dataset construction, has their reliability been evaluated?
Task MAE (Claude) MAE (GPT-4o)
Alert Explanation 0.8 1.0
Retrieved Security Doc QA 0.7 1.1
Suspicious Command Analysis 0.4 1.0
Security Event Query Generation 1.0 0.8
Terraform Security Misconfiguration Fix 1.1 0.4
Average 0.8 0.86
Table 11: Mean absolute error (MAE) between human expert scores and LLM scores across different Primus-Instruct tasks.
Yes, we conducted an experiment to measure the discrepancy between human experts and LLM judges under identical prompts. Specifically, in Sec.2.4 we used Claude 3.5 Sonnet to rate the helpfulness of responses in Primus-Instruct, discarding those that were not helpful enough [^17] The judge prompt is provided in the Appx.E (Fig.11). To validate Claude’s reliability as a judge, we randomly selected ten examples per task for human experts to score, then computed the differences between human, GPT-4o, and Claude ratings.
The discrepancies are reported in Tab.11. Since Primus-Instruct’s responses were generated by GPT-4o, we found that it tended to favor its own answers, which is consistent with findings in LLM-as-a-Judge Zheng et al. (2023). This resulted in slightly larger discrepancies compared to Claude. Based on these results, we found that the gap between LLM-based and human scoring remained within an acceptable range.
• Q4: What is the training objective of Primus-Instruct?
We would like to clarify that our primary goal with the SFT data was not to further improve the model’s cybersecurity capabilities. Instead, our goal was to help the model regain its instruction-following ability without forgetting the cybersecurity knowledge acquired during pre-training. This can be viewed as a continual learning problem involving two tasks: "retaining cybersecurity knowledge" and "learning instruction following". According to LAMOL Sun et al. (2019), language models often suffer from catastrophic forgetting when trained sequentially on multiple tasks—learning a new task tends to overwrite knowledge from previous ones.
A common solution is to interleave data from previous tasks into the new task to mitigate forgetting. Inspired by this, we designed our cybersecurity SFT data to combine both instruction-following and domain-specific knowledge, hoping that the model would learn instruction-following while retaining its earlier cybersecurity understanding. As shown in Tab.6, the results suggest that the model was able to recover instruction-following ability without significant loss in cybersecurity performance.
• Q5: Why does Primus-Instruct appear to have a relatively small number of samples (~1k)?
Task Samples Accepted
Alert Explanation 400 100
Retrieved Security Doc QA 400 100
Suspicious Command Analysis 400 100
Security Event Query Generation 400 100
Terraform Security Misconfiguration Fix 300 96
Total 1,900 496
 + General Instruction Following (339) 2,239 835
Table 12: Initially designed (unfiltered) and accepted (filtered) sample counts per task, where accepted refers to the top 100 samples with a judge score $\geq 8$.
Model Base Model for Merge Merge Model 1 (Task Vector 1) Merge Model 2 (Task Vector 2) Cybersecurity Agg. Score MT-Bench
Llama-Primus-Merged (from unfiltered SFT) Llama-3.1-8b Llama-Primus-Base -> SFT (2,239 samples) Llama-3.1-8b-Instruct 2.44 7.97
Llama-Primus-Merged (from filtered SFT) Llama-3.1-8b Llama-Primus-Base -> SFT (835 samples) Llama-3.1-8b-Instruct 2.63 8.29
Llama-3.1-8b-Instruct 2.29 8.35
Table 13: Comparison of merged Primus models using different versions of the SFT dataset on cybersecurity and MT-Bench benchmarks. The first row refers to applying SFT on Llama-Primus-Base using the unfiltered 2,239 samples from Tab.12 before merging with Llama-3.1-8B-Instruct, while the second row uses the filtered high-quality 835-sample version for SFT prior to merging.
In fact, Primus-Instruct was selected from a larger pool of data. For each task, we initially generated 300–400 samples and rated their helpfulness (on a scale of 1 to 10) using the judge prompt in Fig.11. Only the top 100 samples with scores of at least 8 were retained (Tab.12).
Since we first performed SFT and then merged the resulting model with Llama-3.1-8B-Instruct to balance cybersecurity capabilities and instruction-following ability (Sec.4.3), the SFT and merging steps should be considered as a unified process. We therefore evaluated the combined effect of both. Specifically, we conducted SFT on Llama-Primus-Base separately using both the unfiltered version (2,239 samples) and the filtered high-quality version (835 samples) from Tab.12. Each resulting SFT model was then merged with Llama-3.1-8B-Instruct for comparison.
The merging process involves subtracting each model’s weights from the same base model (Llama-3.1-8B) to obtain two task vectors: one representing cybersecurity knowledge, and the other representing instruction-following ability. The results are shown in Tab.13. We found that applying SFT with a small amount (835) of high-quality data on Llama-Primus-Base before merging yields the best results in both the Cybersecurity Aggregate Score (2.63) and the MT-Bench score (8.29). This is why we chose the filtered high-quality version as Primus-Instruct.
• Q6: Were more baselines compared?
Benchmark ZySec-AI/ SecurityLLM HackMentor/ Llama-7b-lora-iio HackMentor/ Vicuna-7B-lora-iio Llama-Primus- Merged
CISSP 0.6012 0.2908 0.4519 0.7191
CTI-MCQ 0.5676 0.4184 0.5104 0.6656
CTI-RCM 0.4420 0.2770 0.2810 0.6620
CTI-ATE 0.0286 0.2671 0.1411 0.3387
CTI-VSP 1.3923 2.1172 1.6205 1.1233
CyberMetric 0.8140 0.3640 0.6760 0.8660
SecEval 0.4641 0.3640 0.3413 0.5062
Table 14: Performance comparison with existing cybersecurity LLMs across benchmarks. CTI-VSP is scored using Mean Absolute Deviation (lower is better), CTI-ATE uses F1 score, and the others use accuracy. The best results are highlighted in bold.
As shown in Fig.2, most existing cybersecurity-specific LLMs are fine-tuned for narrow tasks, such as password strength detection or malware detection from assembly code. Studies aimed at improving general cybersecurity domain knowledge in LLMs are relatively rare, and to the best of our knowledge, we are the first to pursue this through pre-training.
The primary goal of our comparisons is to demonstrate the effectiveness of our dataset by showing the performance gains of the same base model before and after training on it. Comparisons with other cybersecurity LLMs are difficult to interpret fairly due to differences in training methods and base models. However, to make our findings more convincing, we also identified existing models that incorporate domain knowledge into LLMs via SFT or DPO, and conducted comparisons with them. As shown in Tab.14, our model consistently outperforms these alternatives Zhang et al. (2023a).
Appendix B Primus-Nemotron-CC
We further extracted cybersecurity-related text from Nemotron-CC Su et al. (2024), which claims higher quality and more “unique” tokens than FineWeb (i.e., tokens remaining after global fuzzy deduplication). We scored each Nemotron-CC sample using the binary classifier trained in Sec.2.3 and partitioned the scores into multiple intervals. For each score interval, we sampled 1,000 examples, grouped them by length, sent them to GPT-4o-mini [^18] The prompt is provided in Appx.E (Fig.10) to verify whether they were truly cybersecurity-related, and then calculated the proportion of confirmed samples. The results are shown in Fig.6.
Figure 6: Ratio of cybersecurity-related text across different score bins in Nemotron-CC, grouped by sample length.
Figure 7: Ratio of cybersecurity-related text across score bins in the 1.0 ~ 0.9 range in Nemotron-CC.
Cybersecurity Score Bin Filtered Tokens Dedup.
0.98 ~ 0.85 2.22B 2.05B
0.98 ~ 0.30 4.07B 3.75B
0.98 ~ 0.05 6.02B 5.53B
0.98 ~ 0.0175 8.31B 7.63B
0.98 ~ 0.015 8.89B 8.86B
0.98 ~ 0.01 10.97B 10.05B
0.98 ~ 0.0075 13.10B 11.98B
Table 15: Token counts before and after deduplication for Primus-Nemotron-CC samples (length > 500) across different score bins.
We observed that when sample length is under 500 or the score is below 0.003, the proportion of cybersecurity-related samples falls below 50% in most cases. Therefore, we only retain samples that exceed 500 in length and have a score greater than 0.003. Interestingly, the proportion of cybersecurity samples also declines when the score is very high (> 0.9), likely because our classifier was trained on FineWeb. Thus, we performed a finer-grained analysis on the > 0.9 interval, as shown in Fig.7. Once the score exceeds 0.98, the related proportion drops below 50%, so we only keep samples with scores under 0.98.
Due to computational constraints, we were unable to include all samples that met the above criteria. Instead, we computed the total number of tokens (for samples with length > 500) within different score ranges, as shown in Tab.15. Given our computing budget, we aimed to limit the 70B model’s pretraining dataset to approximately 10B tokens. As a result, we selected the 0.98 ~ 0.175 score range, which contains 7.6B tokens, for inclusion in Primus-Pretraining. This dataset will also be released.
Appendix C CTI-Bench
CTI-Bench is a benchmark for evaluating the reasoning and knowledge capabilities of LLMs in CTI. It consists of several subtasks, including CTI-RCM, CTI-VSP, CTI-ATE, and CTI-MCQ, which assess a model’s ability to analyze vulnerabilities, infer security risks, extract attack techniques, and understand cybersecurity concepts. The following paragraphs present a overview of each subtask.
@@ Appendix C CTI-Bench
Appendix D Training Hyperparameters
D.1 Pre-Training
[8B Model]
Provider: AWS
Framework: NeMo
Hardware: 4 nodes, each with 8 $\times$ H200
Training Time: 30 hours (Primus-Seed+Primus-FineWeb)
Epochs: 2
Learning Rate: 1e-6
Pipeline Model Parallel Size: 4
Tensor Model Parallel Size: 8
Context Parallel Size: 1
Global Batch Size: 12
Micro Batch Size: 12
Warmup Ratio: 0.05
Scheduler: Cosine Annealing
Sequence Length: 16,384
[70B Model]
Provider: NVIDIA
Framework: NeMo
Hardware: 4 nodes, each with 8 $\times$ H100
Training Time: 175 hours
Epochs: 2
Learning Rate: 1e-6
Pipeline Model Parallel Size: 4
Tensor Model Parallel Size: 8
Context Parallel Size: 1
Global Batch Size: 8
Micro Batch Size: 1
Warmup Ratio: 0.05
Scheduler: Cosine Annealing
Sequence Length: 11,264
D.2 Instruction Fine-Tuning
Provider: Azure
Framework: LLaMA-Factory
Hardware: 4 $\times$ A100
Training Time: 2 hours
Epochs: 2
Learning Rate: 1e-6
Deepspeed: ZeRO Stage-3 with CPU Offload
Per Device Train Batch Size: 1
Warmup Ratio: 0.1
Scheduler: Cosine
Cutoff Length: 16,384
D.3 Reasoning Fine-Tuning
Provider: Azure
Framework: LLaMA-Factory
Hardware: 4 $\times$ A100
Training Time: 2.5 hour
Epochs: 3
Learning Rate: 1e-6
Deepspeed: ZeRO Stage-3 with CPU Offload
Per Device Train Batch Size: 3
Warmup Ratio: 0.1
Scheduler: Cosine
Cutoff Length: 8,192
Appendix E Prompts